# Expanding the reach of diffusing wave spectroscopy and tracer bead microrheology

### v1

@@ III.3 Merging two-cell DWS and Echo data
We first generate an interpolating function for $g_{2}^{\mathrm{TC}}(t)-1$ in Mathematica based on the data recorded using a multiple-tau layout of correlation lag times [60]. Mathematica employs piecewise polynomial interpolation (cubic interpolation for smoothness) to construct the interpolating function [57]. This function is then used to calculate the values of $g_{2}^{\mathrm{TC}}(t)-1$ at $t=\tau^{\mathrm{echo}}$ and its multiples in the linear lag time scheme of the echo correlation function. To merge the two datasets, we use the first six echo data points from $t=\tau^{\mathrm{echo}}=0.04$ s to $t=6\tau^{\mathrm{echo}}=0.24$ s. The choice of six is ad hoc, but it is evident that at least a few overlapping data points are necessary to blend the two measurements effectively, while selecting an excessively large number of overlapping points provides little benefit. For these six points $i=1,2,..,6$ we calculate the ratio $\left[g_{2}^{\text{TC}}(i\times\tau_{\text{Echo}})-1\right]/\left[g_{2}^{\text{Echo}}(i\times\tau_{\text{Echo}})-1\right]$. As shown in the inset of Fig. 1 (b), we find that for our (conservative) choice of $\tau^{\mathrm{TC}}\simeq 0.7$s, this range of data points is well described by Eq. (2), and we extract the parameters $A$ and $\tau_{\mathrm{TC}}$ from a fit to the data. Finally, we multiply the full dataset for $g_{2}^{\mathrm{Echo}}(t)-1$ by $A$, divide the dataset $g_{2}^{\mathrm{TC}}(t)-1$ by $e^{-(t/\tau_{\mathrm{TC}})^{2}}$, and truncate it at $t=6\tau^{\mathrm{echo}}=0.24\,\mathrm{s}$. We merge both datasets into one, and the excellent result is shown in Figure 1(c). For comparison with the conventional approach see also Fig. S1.
We also tested a more aggressive setting $\tau^{\mathrm{TC}}\simeq 0.28$s and find equally good agreement as shown in the Supplemental Material, Table S1 and Figures S2 and S3. Our results suggest that as long as $\tau_{\mathrm{TC}}\geq 6\tau_{\mathrm{Echo}}$, corresponding to a second-cell ICF decay of no more than $1/e$ for the last overlapping data point, the parameters $A$ and $\tau_{\mathrm{TC}}$ can be extracted accurately, and the data can be blended and merged smoothly. The statistical noise concerning $A$ and $\tau_{\mathrm{TC}}$ is now reduced from 10% to about 5%, see Table S1. Access to such smaller $\tau_{\mathrm{TC}}$ values implies that a shorter measurement duration can be sufficient to obtain data of comparable quality to that shown in Figure 1. This enhances the suitability of DWS microrheology for monitoring time-evolving processes, such as gelation, a long-standing objective in soft matter research [61, 62]. Previously, limitations in data precision hindered the effectiveness of this approach. The present implementation overcomes these constraints, and while further testing and parameter optimization are beyond the scope of this work, it is conceivable that a time resolution on the order of a minute or two could be achieved, enabling high-quality and stable broadband DWS microrheology data acquisition during relatively fast gelation processes.
@@ III.4.1 Data preparation
In the previous section, we demonstrated how to blend and merge the two-cell and echo measurements without the need for a calibration sample, achieving a smooth and steady transition between the two datasets. We note that the first few echo data points are relatively widely spaced on the time axis, but they are now combined with the two-cell data over the same time range and instead form a fairly dense dataset. Echo data at larger times are densely spaced compared to the multiple-tau channel layout. This outcome is important, as the particle mean square displacement (MSD) we study is typically represented in a logarithmic format, and microrheology routines often rely on analyzing derivatives of the logarithm of the mean squared displacement, Eq. 5. An unevenly spaced distribution of time channels around $\tau_{\text{Echo}}$ can introduce errors and inconsistencies in the data processing.
III.4.2 CONTIN fitting
We use this dataset as input for a fit with CONTIN. If the processing is restricted to short times, smaller than $t=5\times 10^{-4}\,\mathrm{s}$, we consistently obtain an excellent fit to the data. However, if we attempt to fit the entire dataset with CONTIN, the fit does not always converge, and visible deviations between the fit and the data occur. While beyond the scope of this paper, we believe this issue could be addressed by refining the fitting algorithm or by using another unbiased fit to interpolate and smooth the experimental data. Here, our primary objective is to achieve a largely bias-free extrapolation to $t=0$ and smooth data for short times, enabling the application of the inertia correction described in the following section. Without this step, applying the inertia correction using Eq. (8) would be impossible without discarding the high-frequency data of interest or manually optimizing the normalization and smoothing the short-time data as done in earlier work [7, 54]. With our approach, however, the inertia correction is stable and reproducible in all cases we examined. Fig. 2 shows a merged dataset along with the CONTIN fit to the data for $t\leq 5\times 10^{-4}\,\mathrm{s}$. We replace the experimental data with the fit values over the range covered and normalize the data using the value of the fitted curve at $t\to 0$. This modified dataset is then used for further analysis. In the inset of Fig. 2, we show three repetitions of the same experiment shown in the main panel to demonstrate its stability and reproducibility.
Figure 2: CONTIN fit. Red circles: Merged correlation function $g_{2}(t)-1$ as shown in Figure 1(c). Blueline: CONTIN-regularized fit to the correlation function for times $t\leq 5\times 10^{-4}\,\mathrm{s}$. Both curves are normalized using the CONTIN fit extrapolated to $t=0$. Inset: Merged $g_{2}(t)-1$ from three repetitions of the experiment at $T=20^{\circ}\mathrm{C}$.
@@ III.4.3 Microrheology
We determine the samples mean free path to $\ell^{\ast}\simeq 325\mu$m using the commercial DWS RheoLab instrument (LS Instruments, Switzerland) along with the known values of $L=2\,\mathrm{mm}$ and $k=\frac{2\pi n}{\lambda}=13.2\,\mu\mathrm{m}^{-1}$, for $n_{\mathrm{water}}=1.33$, see also Figure S4. From this, we extract the particle mean squared displacement $\langle\Delta r^{2}(t)\rangle$ using Eq. (4) and the data is shown in Fig. 3. To smooth the data for times where the CONTIN fit was not applied, we use a sixth-order polynomial fit, shown in Fig. 3, as is commonly done [25], and replace the data for $t\geq 5\times 10^{-4}$s with the fitted values. We note once again that using a polynomial fit at very short times is not advisable, as the mean squared displacement exhibits a very complex time dependence due to the effects of inertia.
Figure 3: Open red circles: Natural Logarithm of the mean squared particle displacement (MSD) calculated from the intensity correlation function via Eq. (4). Solid line: Polynomial fit of order six to the data displayed for times $t\geq 3\times 10^{-7}\,\mathrm{s}$.
With knowledge of the bead radius ($a=210\,\mathrm{nm}$), we determine the apparent complex modulus $Z^{\ast}(\omega)$ using Eq. (7) and the result is shown in Fig. 4 (a). Finally we apply Eq. (8) to account for the fluid and bead inertia, obtaining the true moduli of the sample, shwon in Fig. 4 (b). The following values are used: $\rho_{b}=1050\,\mathrm{kg/m^{3}}$ (polystyrene bead mass density) and $\rho\simeq 1000\,\mathrm{kg/m^{3}}$ (medium density). The water viscosity is $\eta_{s}=1.02$mPas at T$=20^{\circ}$.
Figure 4: (a) Apparent complex moduli of the micellar solution $Z^{\ast}(\omega)$ at T$=20^{\circ}$, as well as the real and the imaginary part $Z^{\prime}(\omega),Z^{\prime\prime}(\omega)$, obtained using the Generalized Stokes-Einstein Relation (GSE), Eq. (7), and Mason’s routine to convert the mean squared displacement [50]. (b) Corrected complex modulus of the sample after accounting for bead and fluid inertia, which primarily affects data at frequencies $\omega\geq 10^{5}\,\mathrm{rad/s}$. The solvent loss modulus $G^{\prime\prime}\left(\omega\right)=\eta_{s}\omega$ is shown as a dashed line. The data is truncated at $\omega=5\times 10^{6}\,\mathrm{rad/s}$.
In the high-frequency limit, we truncated the data at $\omega\simeq 5\times 10^{6}\,\mathrm{rad/s}$, above which the ICF decays by less 0.5% from 1 to 0.995. Using larger $L/\ell^{\ast}$ leads to a faster decay and may provide access to even higher frequencies. However, as seen in Fig. 4 (b), the complex modulus gradually approaches the modulus of the solvent $G^{\prime\prime}\left(\omega\right)=\eta_{s}\omega$. Thus, the upper bound for a stable DWS high-frequency analysis remains to be determined and is likely system dependent. Using the improved DWS implementation, we are already pushing the boundaries of the technique. Still, both the blending and merging procedure as well as the inertia correction are stable, and the process is highly reproducible.
@@ IV Conclusions
In the present work, we presented a new implementation of two-cell echo diffusing wave spectroscopy (DWS) aimed at significantly improving the data quality and stability and thus extending the reach of DWS and DWS-based tracer bead microrheology. We replaced the standard calibration method for two-cell DWS, introduced by Romer et al. in 2000 [36], later used by many others and implemented in commercial DWS equipment [41, 29, 13], with a calibration-free approach. The new method generically corrects small fluctuations in the terminal decay time arising from imperfect temporal averaging, allowing for a seamless transition between two-cell DWS and echo DWS. This smooth transition, in turn, facilitates a stable conversion of the bead mean square displacement to the complex viscoelastic modulus via the generalized Stokes-Einstein relation (GSER). Combined with the unbiased normalization of the correlation function using the CONTIN fit and the elimination of shot noise, the quality of DWS data is substantially improved. This enhancement enables the routine application of published formulas for correcting the effects of fluid and tracer bead inertia. As a result, the method provides unprecedented access to the high-frequency regime [7]. This capability will be beneficial for addressing questions related to the short-time dynamics of polymeric and other soft matter systems. Furthermore, this approach brings DWS closer to advanced techniques, such as neutron spin echo spectroscopy, connecting the high frequency viscoelastic relaxations to sub-microsecond scale dynamics in systems like microgels [63] or protein hydrogels [64].
@@ Acknowledgements.
We thank Andrea Vaccaro (LS Instruments, Switzerland) for valuable discussions and assistance with the CONTIN fit algorithm. This work was financially supported by the Swiss National Science Foundation through project grant no. 10000141.
@@
F.S. designed the study and wrote the Mathematica code. M.H. and C.Z. conducted the experiments. F.S. performed the data analysis with contributions from C.Z. and M.H. F.S. drafted the manuscript.
@@
All experimental data and model output discussed in the manuscript will be uploaded to the repository Zenodo (xxxxxxxx). All additional data sets generated during and/or analyzed during the current study are available from the corresponding author upon reasonable request.
Supplemental Material
Supplementary Figures
Figure S1: Two-cell DWS using the conventional calibration with a solid reference sample (Teflon). The data corresponds to that shown in Figure 1 of the main text, taken at $T=20^{\circ}$. Here, we divided the recorded $g_{2}^{\text{TC}}(t)-1$ by the ICF of the solid calibration sample, $g_{2}^{\text{2nd}}(t)-1$, shown as a dashed line in Figure 1 (a). The resulting $\big{(}g_{2}(t)-1\big{)}$ for three repetitions is shown and plotted up to $t=\tau_{\text{TC}}$. Merging the data with $\big{(}g_{2}^{\text{Echo}}(t)-1\big{)}$ may lead to an unsteady curve with a kink in many cases, most prominently in the present example for the blue line.
Figure S2: Same experiment as in Figure 1, but with 2.5 times faster motion of the diffusor during the two-cell measurement, leading to $\tau_{\text{TC}}\simeq 0.28\ \mathrm{s}$. The value of $\ell^{\ast}\simeq 413\ \mu\mathrm{m}$ is slightly higher due to minor settling of the tracer beads. In the overlap region, indicated by the dashed lines in panel (a), where the two datasets are blended and merged, the TC-ICF now decays to about one-half, $e^{-\left(6\tau_{\text{Echo}}/0.28\ \mathrm{s}\right)^{2}}\simeq 1/2$. The fit with Eq. (2) remains excellent and stable, and the merged ICF is continuous and smooth, as shown in panels (b) and (c).
Figure S3: Five repetitions of the experiment shown in Fig. S1. The measurement times are 300 seconds for the TC-DWS and 60 seconds for the Echo-DWS. The fitted parameters $A$ and $\tau_{\text{TC}}$ are listed in Table S1.
Experiment No. Amplitude $A$ Decay Time $\tau_{\text{TC}}$ [s]
1 0.850782 0.287184
2 0.763751 0.286332
3 0.752708 0.287966
4 0.794214 0.275153
5 0.755180 0.277028
Table S1: Summary of experimental results showing the fitted amplitude and decay time for each of the experiments shown in Fig. S3. The mean values are $\langle A\rangle=0.7833\pm 0.0368$ and $\langle\tau_{\text{TC}}\rangle=0.2827\pm 0.0055$.
Figure S4: Comparison of the $T=25^{\circ}$ data obtained using a commercial DWS instrument (red symbols), DWS RheoLab (LS Instruments, Switzerland), and the data obtained with our DWS microrheology implementation, as shown in Fig. 5 of the manuscript (blue symbols). The exact same sample contained in a $L=2$ mm cuvette was analyzed. Storage modulus $G^{\prime}(\omega)$ and loss modulus $G^{\prime\prime}(\omega)$ are shown as full and open symbols, respectively.

### v2

@@ III.3 Merging two-cell DWS and Echo data
We first generate an interpolation function for $g_{2}^{\mathrm{TC}}(t)-1$ with SciPy based on the data recorded using a multiple tau layout of correlation lag times [72]. SciPy employs piecewise polynomial interpolation (cubic interpolation for smoothness) to construct the interpolating function [73]. This function is then used to calculate the values of $g_{2}^{\mathrm{TC}}(t)-1$ at $t=\tau^{\mathrm{echo}}$ and its multiples in the linear lag time scheme of the echo correlation function. To merge the two datasets, we use the first six echo data points from $t=\tau^{\mathrm{echo}}=0.04$ s to $t=6\tau^{\mathrm{echo}}=0.24$ s. The choice of six is ad hoc, but it is evident that at least a few overlapping data points are necessary to blend the two measurements effectively, while selecting an excessively large number of overlapping points provides little benefit. For these six points $i=1,2,..,6$ we calculate the ratio $\left[g_{2}^{\text{TC}}(i\times\tau_{\text{Echo}})-1\right]/\left[g_{2}^{\text{Echo}}(i\times\tau_{\text{Echo}})-1\right]$. As shown in the inset of Fig. 1 (b), we find that for $\tau^{\mathrm{Echo}}\simeq 0.04$s and our choice of $\tau^{\mathrm{TC}}\simeq 0.29$s, this range of data points is well described by Eq. (2), and we extract the parameters $A$ and $\tau_{\mathrm{TC}}$ from a fit to the data. Finally, we multiply the full dataset for $g_{2}^{\mathrm{Echo}}(t)-1$ by $A$, divide the dataset $g_{2}^{\mathrm{TC}}(t)-1$ by $e^{-(t/\tau_{\mathrm{TC}})^{2}}$, and truncate it at $t=6\tau^{\mathrm{echo}}=0.24\,\mathrm{s}$. We merge both data sets into one, and the result is shown as a red line in Figure 1(c). Our results suggest that as long as $\tau_{\mathrm{TC}}\geq 6\tau_{\mathrm{Echo}}$, corresponding to a second-cell ICF decay of no more than $1/e$ for the last overlapped data point, the parameters $A$ and $\tau_{\mathrm{TC}}$ can be extracted accurately and the data can be seamlessly blended and merged. We note that the blending and merging approach introduced here could equally be applied when combining two-cell DWS with camera-based DWS to cover the slow relaxation time window, as discussed in Refs. [40, 41].
@@ III.4.1 Data preparation
In the previous section, we demonstrated how to blend and merge two-cell and echo measurements without the need for a calibration sample, achieving a smooth and steady transition between the two datasets. We note that the first few echo data points are relatively widely spaced on the time axis, but they are now combined with the two-cell data over the same time range and instead form a fairly dense data set. Echo data at larger times are densely spaced compared to the multiple-tau channel layout. This outcome is important, as the mean square displacement (MSD) of the particles is typically represented in logarithmic format, and microrheology routines often rely on analyzing derivatives of the logarithm of the mean squared displacement, Eq. (10). An unevenly spaced distribution of time channels around $\tau_{\text{Echo}}$ can introduce errors and inconsistencies in data processing.
III.4.2 Fitting and smoothing
We use the merged data set as input for the exponential basis fit, Section II.2.1. Our primary objective is to achieve a minimally biased extrapolation to $t=0$ and smooth data, enabling the application of the inertia correction and the microrheology algorithm. Using the SciPy interpolation function, we remap the data onto the original multiple-tau channel layout of the digital correlator to prevent oversampling and undue weighting of the longer-time data. The merged data set and the fit $g_{2}^{\text{fit}}(t)-1$ are shown in Figure 2 (a) together with the residuals, Figure 2 (b). In the inset of Figure 2 a), we show five repetitions of the same experiment to demonstrate its stability and reproducibility, see also supplementary Table 1. We note that without this step, applying the inertia correction using Eq. (13) would not be feasible without either discarding the high-frequency data of interest or manually adjusting the normalization and smoothing the short-time data, as done in earlier work [7, 65].
Figure 2: (a) Fit to the intensity correlation function (ICF). Red circles: Merged correlation function $g_{2}(t)-1$ as shown in Figure 1(c). Blue line: Exponential basis fit $g_{2}^{\text{fit}}(t)$ to the correlation function. The experimental data is normalized using the ICF fit extrapolated to $t=0$. Inset: Five repetitions of the merged correlation function $g_{2}(t)-1$, normalized using the exponential basis fit. (b) Open circles: Fit residuals $r(t)=g_{2}(t)-g_{2}^{\text{fit}}(t)$. (c) Influence of the number of exponentials on the fit quality. Open circles: global fit error, Eq. (7), evaluated for different numbers of exponential terms NN.
@@ III.4.3 Microrheology
[…]
Figure 3: Open black circles: Mean squared particle displacement (MSD) calculated from the intensity correlation function via Eq. (9).
With knowledge of the radius of the beads ($a=210\,\mathrm{nm}$), we determine the apparent complex modulus $Z^{\ast}(\omega)$ using Eq. (12) and the result is shown in Fig. 4 (a). Finally, we apply Eq. (13) to account for the fluid and bead inertia, obtaining the true moduli of the sample, shown in Fig. 4 (b), see also Figure S2. The following values are used: $\rho_{b}=1050\,\mathrm{kg/m^{3}}$ (polystyrene bead mass density) and $\rho\simeq 1000\,\mathrm{kg/m^{3}}$ (medium density) and temperature-dependent water viscosity $\eta_{s}$ [74].
[…]
In the high-frequency limit, we truncated the data at $\omega\simeq 5\times 10^{6}\,\mathrm{rad/s}$, above which the ICF decays by less 0.5% from 1 to 0.995. Using a larger $L/\ell^{\ast}$ leads to a faster decay and may provide access to even higher frequencies. However, as seen in Fig. 4 (b), the complex modulus gradually approaches the modulus of the solvent $G^{\prime\prime}\left(\omega\right)=\eta_{s}\omega$. Thus, the upper bound for a stable DWS high-frequency analysis remains to be determined and is likely to be system dependent. Using the improved DWS implementation, we are already pushing the boundaries of the technique. Still, both the blending and merging procedures and the inertia correction are stable, and the process is highly reproducible.
@@
Finally, we also demonstrated the method on a viscoelastic solid sample (Figure S5 and S6). To this end, we studied a dense microgel suspension without added tracer beads. The system is identical to the one described in  [75].
@@ IV Conclusions
In the present work, we presented a new implementation of two-cell echo diffusing wave spectroscopy (DWS) aimed at significantly improving the data quality and stability and thus extending the reach of DWS and DWS-based tracer bead microrheology. We replaced the standard calibration method for two-cell DWS, introduced by Romer et al. in 2000 [36], later used by many others and implemented in commercial DWS equipment [42, 29, 13], with a calibration-free approach. The new method generically corrects small fluctuations in the terminal decay time arising from imperfect temporal averaging, allowing for a seamless transition between two-cell DWS and echo DWS. This smooth transition, in turn, facilitates a stable conversion of the bead mean square displacement to the complex viscoelastic modulus via the generalized Stokes-Einstein relation (GSER). Combined with the normalization of the correlation function using the exponential basis fit and the elimination of shot noise, the quality of DWS data is substantially improved. This enhancement enables the routine application of published formulas to correct the effects of fluid and tracer bead inertia. As a result, the method provides unprecedented access to the high-frequency regime [7]. This capability will be beneficial for addressing questions related to the short-time dynamics of polymeric and other soft-matter systems. Furthermore, this approach brings DWS closer to other advanced techniques, such as neutron spin echo spectroscopy, connecting the high-frequency viscoelastic relaxations to sub-microsecond scale dynamics in systems like microgels [76] or protein hydrogels [77].
@@ Acknowledgements.
We thank Erika Eiser (NTU Trondheim, Norway) and Andrea Vaccaro (LS Instruments, Switzerland) for valuable discussions. This work was financially supported by the Swiss National Science Foundation through project grant #10000141 and #10002524. We thank the referees for their constructive feedback, and in particular the second referee for helpful suggestions on improving the fitting of the ICF.
@@
F.S. designed the study and wrote the original Mathematica code. M.H. and C.Z. conducted the experiments. C.Z. designed the exponential basis fit. M.H. implemented and tested the Python package (pyDWS). All authors participated in the data analysis. F.S. wrote the manuscript with contributions from all authors.
@@
All experimental data and model output discussed in the manuscript have been uploaded to the repository Zenodo (17426137). The source code for the pyDWS Python package can be found on GitHub/pyDWS and can also be installed directly from PyPi/pyDWS. All additional data sets generated and / or analyzed during the current study are available from the corresponding author upon reasonable request.